OpenStudy (vegan555):

If Ben invests $4000 at 4% interest per year, how much additional money must he invest at 5 1/2% annual interest to ensure that the interest he receives each year is 4 1/2% of the total amount invested?

OpenStudy (anonymous):

Let "x" be the amount of additional money he needs to invest. EQUATION: interest + interest = interest. 0.04(4000) + 0.055x = 0.045(4000+x) 160 +0.055x = 180+0.045x 0.010x=20 x= $2000 (amount of additional money he needs to invest at 5 1/2 %) Cheers.

OpenStudy (vegan555):

thank u!!

OpenStudy (anonymous):

You're Welcome! :)